OpenStudy (anonymous):

Solve for x: (1/2)lnx-ln(3x^2 +2) = ln (1/5)

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ (1/2)\ln x-\ln(3x^2 +2) = \ln (1/5)\LARGE\color{white}{ \rm │ }}\) like this ?

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ (1/2)\ln x-\ln(3x^2 +2) = \ln (1/5)\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ \ln \sqrt{~x}~ -\ln(3x^2 +2) = \ln (1/5)\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ \ln[~~\sqrt{~x}/(3x^2 +2)~~] = \ln (1/5)\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ \sqrt{~x}/(3x^2 +2)~~] = 1/5\LARGE\color{white}{ \rm │ }}\)

OpenStudy (solomonzelman):

now, cross multiply, \(\normalsize\color{blue}{ \sqrt{~x}/(3x^2 +2)= 1/5\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ 5\sqrt{~x}=3x^2 +2\LARGE\color{white}{ \rm │ }}\)

OpenStudy (solomonzelman):

Then raise both sides to the second power, \(\normalsize\color{blue}{ 25x=9x^4 +12x^2+4\LARGE\color{white}{ \rm │ }}\)

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{9x^4 +12x^2-25x+4)=0\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ (x-1)(9x^3 +9x^2+21x-4)=0\LARGE\color{white}{ \rm │ }}\)

OpenStudy (solomonzelman):

x-1=0, x=1

OpenStudy (solomonzelman):

Solve for x from \(\normalsize\color{blue}{9x^3 +9x^2+21x-4=0\LARGE\color{white}{ \rm │ }}\)